# Lighter, Better, Faster Multi-Source Domain Adaptation with Gaussian Mixture Models and Optimal Transport

### v1

@@ Abstract
Keywords:
Domain Adaptation Optimal Transport Gaussian Mixture Models.
@@ 1 Introduction
Summary of Contributions. Our contributions are threefold: 1. We propose a novel strategy for mapping the parameters of GMMs using OT (section 3.1 and theorem 3.1); 2. We propose a novel algorithm for computing mixture-Wasserstein barycenters of GMMs (algorithm 1 in section 3.3); 3. We propose an efficient parametric extension of the WBT and DaDiL algorithms based on GMMs (section 3.4). We highlight that, while GMMs were previously employed in single source DA [8], to the best of our knowledge this is the first work to leverage GMM-OT for MSDA.
@@ 2 Preliminaries
From a theoretical standpoint, this framework is useful because $\mathcal{R}_{Q}$ is bounded by ^$\hat{\mathcal{R}}_{Q}$ and a complexity term depending on the number of samples $n$, and the Vapnik-Chervonenkis dimension of $\mathcal{H}$ [28, Section 6]. As a result, $\hat{h}$ minimizing the empirical risk is guaranteed to generalize to unseen samples of $P$.
@@
Optimal Transport is a field of mathematics concerned with the displacement of mass at least effort [21, 15]. Given probability measures $P,Q\in\mathbb{P}(\mathcal{X})$, the Monge formulation [21, Section 2.2.] of OT seeks for a mapping $T$,
@@
where $T_{\sharp}$ is the push-forward mapping of $T$, i.e., $T_{\sharp}P(A)=P(T^{-1}(A))$. Nonetheless, this problem poses technical difficulties, mainly due the constraint $T_{\sharp}P=Q$. A more tractable formulation was proposed by Kantorovich [21, Section 2.3.], and relies on OT plans,
@@ 3.1 First Order Analysis of $\mathcal{MW}_{2}$
Equation 10 is similar to the barycentric mapping in Empirical Optimal Transport (EOT) [2, eq. 13], which serves as an approximation for the Monge mapping between $P$ and $Q$. In our case, the barycentric mappings act on the parameters of the GMM, rather than on its samples. Theorem 3.1 will be useful in the calculation of $\mathcal{MW}_{2}$ barycenters.
@@ 3.2 Supervised Mixture-Wasserstein Distances
Theorem 3.2
Under the same conditions of theorem 3.1, let $P_{i}$ and $Q_{j}$ be equipped with labels $\mathbf{y}_{i}^{(P)}$ and $\mathbf{y}_{j}^{(Q)}$. The first order optimality conditions of $\mathcal{SMW}_{2}$ with respect $\mathbf{m}_{i}$ and $\mathbf{s}_{i}$ are given by eq. 10. Furthermore, for $\mathbf{y}_{i}$,
@@ 3.4 Multi-Source Domain Adaptation through GMM-OT
GMM-WBT. The intuition of this algorithm is transforming the MSDA scenario into a single-source one, by first calculating a Wasserstein barycenter of $B=\mathcal{B}(\mathds{1}_{N}/N;\mathcal{Q}_{S})$. After this step, WBT solves a single-source problem between $B$ and $Q_{T}$. When each $Q_{S_{\ell}}$ is a GMM, the parameters of $B$ are estimated through algorithm 1. Next, one solves for $\omega^{(T)}=\text{GMMOT}(B,Q_{T})$, so that the parameters of $B$ are transported towards $Q_{T}$ using theorems 3.1 and 3.2,

### v2

@@ Abstract
Keywords Domain Adaptation  $\cdot$ Optimal Transport  $\cdot$ Gaussian Mixture Models.
@@ 1 Introduction
Summary of Contributions. Our contributions are threefold: 1. We propose a novel strategy for mapping the parameters of GMMs using OT (section 3.1 and theorem 1); 2. We propose a novel algorithm for computing mixture-Wasserstein barycenters of GMMs (algorithm 1 in section 3.3); 3. We propose an efficient parametric extension of the WBT and DaDiL algorithms based on GMMs (section 3.4). We highlight that, while GMMs were previously employed in single source DA [14, 15], to the best of our knowledge this is the first work to leverage GMM-OT for MSDA.
@@ 2 Preliminaries
From a theoretical standpoint, this framework is useful because $\mathcal{R}_{Q}$ is bounded by ^$\hat{\mathcal{R}}_{Q}$ and a complexity term depending on the number of samples $n$, and the Vapnik-Chervonenkis dimension of $\mathcal{H}$ [17, Section 6]. As a result, $\hat{h}$ minimizing the empirical risk is guaranteed to generalize to unseen samples of $P$.
@@
Optimal Transport is a field of mathematics concerned with the displacement of mass at least effort [18, 19]. Given probability measures $P,Q\in\mathbb{P}(\mathcal{X})$, the Monge formulation [18, Section 2.2.] of OT seeks for a mapping $T$,
@@
where $T_{\sharp}$ is the push-forward mapping of $T$, i.e., $T_{\sharp}P(A)=P(T^{-1}(A))$. Nonetheless, this problem poses technical difficulties, mainly due the constraint $T_{\sharp}P=Q$. A more tractable formulation was proposed by Kantorovich [18, Section 2.3.], and relies on OT plans,
@@ 3.1 First Order Analysis of $\mathcal{MW}_{2}$
Equation 10 is similar to the barycentric mapping in Empirical Optimal Transport (EOT) [6, eq. 13], which serves as an approximation for the Monge mapping between $P$ and $Q$. In our case, the barycentric mappings act on the parameters of the GMM, rather than on its samples. Theorem 1 will be useful in the calculation of $\mathcal{MW}_{2}$ barycenters.
@@ 3.2 Supervised Mixture-Wasserstein Distances
Theorem 2.
Under the same conditions of theorem 1, let $P_{i}$ and $Q_{j}$ be equipped with labels $\mathbf{y}_{i}^{(P)}$ and $\mathbf{y}_{j}^{(Q)}$. The first order optimality conditions of $\mathcal{SMW}_{2}$ with respect $\mathbf{m}_{i}$ and $\mathbf{s}_{i}$ are given by eq. 10. Furthermore, for $\mathbf{y}_{i}$,
@@ 3.4 Multi-Source Domain Adaptation through GMM-OT
GMM-WBT. The intuition of this algorithm is transforming the MSDA scenario into a single-source one, by first calculating a Wasserstein barycenter of $B=\mathcal{B}(\mathds{1}_{N}/N;\mathcal{Q}_{S})$. After this step, WBT solves a single-source problem between $B$ and $Q_{T}$. When each $Q_{S_{\ell}}$ is a GMM, the parameters of $B$ are estimated through algorithm 1. Next, one solves for $\omega^{(T)}=\text{GMMOT}(B,Q_{T})$, so that the parameters of $B$ are transported towards $Q_{T}$ using theorems 1 and 2,
@@ 5 Conclusion
Appendix A Introduction
In this supplementary materials, we provide additional details about our theoretical and experimental results.
Appendix B Proofs
B.1 Proof for Theorem 1
Theorem 3.
$\mathbf{m}_{i}=T_{\omega}(\mathbf{m}_{i}^{(P)})=\sum_{j=1}^{K_{Q}}\dfrac{\omega_{ij}^{\star}}{p_{i}}\mathbf{m}_{j}^{(Q)}\text{, and }\mathbf{s}_{i}=T_{\omega}(\mathbf{s}_{i}^{(P)})=\sum_{j=1}^{K_{Q}}\dfrac{\omega_{ij}^{\star}}{p_{i}}\mathbf{s}_{j}^{(Q)}.$ (17)
Proof.
Under the GMM-OT framework, the $2-$mixture Wasserstein distance is given by,
$\mathcal{MW}_{2}(P,Q)^{2}=\sum_{i=1}^{K_{P}}\sum_{j=1}^{K_{Q}}\omega_{ij}^{\star}\mathcal{W}_{2}(P_{i},Q_{j})^{2},$
where $\omega_{ij}^{\star}$ is the mass transported from component $i$ in $P$ to component $j$ in $Q$ by the solution to the GMM-OT problem. Since $P_{i}$ and $Q_{j}$ are axis-aligned Gaussian measures,
$\mathcal{W}_{2}(P_{i},Q_{j})^{2}=\lVert\mathbf{m}_{i}-\mathbf{m}_{j}^{(Q)}\rVert_{2}^{2}+\lVert\mathbf{s}_{i}-\mathbf{s}_{j}^{(Q)}\rVert_{2}^{2},$
where $\mathbf{m}_{i}$ and $\mathbf{s}_{i}$ are the free parameters we want to determine. Henceforth we derive the first-order conditions for $\mathbf{m}_{i}$. The derivations for $\mathbf{s}_{i}$ are completely analogous. By linearity,
$\dfrac{\partial\mathcal{MW}_{2}^{2}}{\partial\mathbf{m}_{i}}=2\sum_{j=1}^{K_{Q}}\omega_{ij}^{\star}(\mathbf{m}_{i}-\mathbf{m}_{j}^{(Q)})=2\biggr{(}p_{i}\mathbf{m}_{i}-\sum_{j=1}^{K_{Q}}\omega_{ij}^{\star}\mathbf{m}_{j}^{(Q)}\biggr{)},$
equating this last term to 00, one gets the desired equality. ∎
The result of theorem 3 is quite similar to the barycentric mapping of [33, 6] which was previously used to map points in $P$ to points in $Q$. Here, we transport the parameters of GMMs instead.
B.2 Proof of Theorem 2
Theorem 4.
Under the same conditions of theorem 3, let $P_{i}$ and $Q_{j}$ be equipped with labels $\mathbf{y}_{i}^{(P)}$ and $\mathbf{y}_{j}^{(Q)}$. The first order optimality conditions of $\mathcal{SMW}_{2}$ with respect $\mathbf{m}_{i}$ and $\mathbf{s}_{i}$ are given by eq. 17. Furthermore, for $\mathbf{y}_{i}$,
$\mathbf{y}_{i}=T_{\omega}(\mathbf{y}_{i}^{(P)})=\sum_{j=1}^{K_{Q}}\dfrac{\omega_{ij}^{\star}}{p_{i}}\mathbf{y}_{j}^{(Q)}.$ (18)
Proof.
Note that the label distance term added to $\mathcal{SMW}$ does not interfere with the terms on $\mathbf{m}_{i}$ and $\mathbf{s}_{i}$, hence the optimality conditions with respect these variables remain unchanged. Due the independence of these terms, the first-order optimality condition with respect $\mathbf{y}_{i}$ is,
$\dfrac{\partial\mathcal{SMW}_{2}^{2}}{\partial\mathbf{y}_{i}}=2\beta\sum_{j=1}^{K_{Q}}\omega_{ij}^{\star}(\mathbf{y}_{i}-\mathbf{y}_{j}^{(Q)})=2\beta\biggr{(}p_{i}\mathbf{y}_{i}-\sum_{j=1}^{K_{Q}}\omega_{ij}^{\star}\mathbf{y}_{j}^{(Q)}\biggr{)},$
which, for $\beta>0$, is zero if and only if $\mathbf{y}_{i}=T_{\omega}(\mathbf{y}_{i}^{(P)})$. ∎
B.3 Derivation of Algorithm 1
In this section, we derive the first-order conditions for the functional $B\mapsto\sum_{c}\lambda_{c}\mathcal{SMW}_{2}(B,P_{c})^{2}$. Let $B$ be parametrized by $\theta_{B}=\{(\mathbf{m}_{i}^{(B)},\mathbf{s}_{i}^{(B)},\mathbf{y}_{i}^{(B)})\}_{i=1}^{K_{B}}$, where these variables are initialized randomly as in algorithm 1. For fixed $\theta_{B}$, we find $\omega_{1}^{\star},\cdots,\omega_{C}^{\star}$ transport plans. Then, for fixed transport plans, we solve,
$\underset{\theta_{B}}{\text{argmin}}\,\mathcal{L}(\theta_{B})$ $=\sum_{c=1}^{C}\lambda_{c}\sum_{i=1}^{K_{B}}\sum_{j=1}^{K_{P}}\omega_{c,i,j}^{(\star)}M_{c,i,j},$
where $\text{where }M_{c,i,j}$ $=\lVert\mathbf{m}_{i}^{(B)}-\mathbf{m}_{j}^{(P_{c})}\rVert_{2}^{2}+\lVert\mathbf{s}_{i}^{(B)}-\mathbf{s}_{j}^{(P_{c})}\rVert_{2}^{2}+\beta\lVert\mathbf{y}_{i}^{(B)}-\mathbf{y}_{j}^{(P_{c})}\rVert_{2}^{2}$
which can be optimized by taking derivatives with respect $\mathbf{m}^{(B)}_{i}$, $\mathbf{s}^{(B)}_{i}$ and $\mathbf{y}^{(B)}_{i}$. For instance, taking the derivative of $\mathcal{L}(\theta_{B})$ with respect $\mathbf{m}_{i}^{(B)}$,
$\dfrac{\partial\mathcal{L}}{\partial\mathbf{m}_{i}^{(B)}}=2\sum_{c=1}^{C}\lambda_{k}\sum_{j=1}^{K_{P}}\omega_{c,i,j}^{\star}(\mathbf{m}_{i}^{(B)}-\mathbf{m}_{j}^{(P_{c})})=\dfrac{2}{K_{B}}\mathbf{m}_{i}^{(B)}-2\sum_{c=1}^{C}\lambda_{c}\sum_{j=1}^{K_{P}}\omega_{c,i,j}^{\star}\mathbf{m}_{j}^{(P_{c})}$
setting the derivative to 00, one has,
$\mathbf{m}_{i}^{(B)}$ $=\sum_{c=1}^{C}\lambda_{c}T_{\omega_{c}^{\star}}(\mathbf{m}_{i}^{(B)}).$
Similar results can be acquired for $\mathbf{s}_{i}^{(B)}$ and $\mathbf{y}_{i}^{(B)}$ by taking the appropriate derivatives.
Appendix C Additional Experiments
Figure 8: Evolution of barycentric coordinates over the simplex $\Delta_{C}$.
Toy Example. We analyze the evolution of the atoms in GMM-DaDiL as training progresses (figure 9). Note that, initially, the means are close to each other, so there are no distinction between classes. Likewise, in our initialization, the labels are uniform histograms over the classes, so that the atoms are not certain about which class they represent. With 50 iterations, DaDiL manages to spread the class clusters, but, at this point, the barycentric coordinates of each domain are still too similar (i.e., $\lambda_{\ell}\approx\mathds{1}_{K}/K$, see figure 2.a in the main paper). As the training progresses, the barycentric coordinates of the different domains begins to spread over the simplex $\Delta_{C}$ (see figure 8). As a result, the atoms begin to be different from each other (compare figures 9 (b) and (c)).
Hyper-parameter Setting: Here we detail the hyper-parameter setting for our method. GMM-WBT involves a single hyper-parameter: the number of components $K$ in the GMMs. DaDiL involves a few other parameters: the number of atoms $C$, the learning rate $\eta$ and the number of optimization steps $N_{iter}$. As [12], we use Adam [34] for optimizing eq. 16 (see main paper). Over all our experiments we set $\eta=10^{-1}$. We show our hyper-parameters in table 3. We remark that, as we analyze in Fig 6. b (main paper), our proposed methods, especially GMM-DaDiL, are robust to the choice of hyper-parameters.
Table 3: Hyper-parameter setting
Benchmark $K$ (GMM-WBT) $K$ (GMM-DaDiL) $C$ $N_{iter}$
Office 31 155 186 3 150
Office-Home 910 455 4 150
TEP 261 174 7 150
CWRU 30 90 5 200
(a) 1st iteration.
(b) 50th iteration.
(c) 100th iteration.
(d) 150th iteration.
(e) 200th iteration.
Figure 9: Evolution of atoms as the training progresses.